package of ballons contains 5 green, 3 yellow, 4 red, and 8 pink balloons. suppose you reach in the package and choose one balloon at random. What is the probability of each event. Express each answer as a fraction, percent, and decimal.
P(RED BALLOON)
P(YELLOW BALLOON)
P(PINK BALLOON)
P(ORANGE BALLOON)
P(RED OR YELLOW BALLOON)
P(NOT GREEN BALLOON)

Answer by ikleyn(53748)

.
package of ballons contains 5 green, 3 yellow, 4 red, and 8 pink balloons.
suppose you reach in the package and choose one balloon at random. What is the probability of each event.
Express each answer as a fraction, percent, and decimal.

P(RED BALLOON)             - 5 of (5+3+4+8) = 5 of 20 = 5/20 = 1/4 = 0.25 = 25%

P(YELLOW BALLOON)          - 3                  of 20 = 3/20 = 0.15 = 15%

P(PINK BALLOON)            - 8                  of 20 = 8/20 = 0.40 = 40%

P(ORANGE BALLOON)                do it on your own

P(RED OR YELLOW BALLOON)   - (4+3)              of 20 = 7/20 = 0.35 = 35%

P(NOT GREEN BALLOON)       - (3+4+8)            of 20 = 15/20 = 3/4 = 0.75 = 75%